A contest offers the winner the following prize. S8 975.00 to be paid in one year with each future payment 5.00% greater than the last. The payoff will go forever! A person considering the contest expects to earn 10.00% per year on his money. What is the value of this prize in todays dollars? Answer Format: Currency: Round to: 2 decimal places. Enter Answer Here Submit Answer Prev Problem All Problems Next Problem
Answer #1

P0 = P1/Ke-G

where P0 is value of price today

P1 is value of price at year 1

Ke is the expectation that a person is expecting

G is the annual growth rate

P0 = P0 ( 1+G)/Ke - G

= 8975 ( 1+0.05)/10%-5%

= 9423.75/5% i.e 188475